The sum of jack n jill are 78,Jack is twice as old as jill.what are their ages?
Answer by addingup(3677)

Let's call Jack "a" and Jill "b".
The problem says that a (Jack) is 2 times b (Jill). Write this in equation form:
2b + b = 78
3b = 78 Divide both sides by 3
b = 26 Great! Jill is 26, and Jack is 26*2 = 52
Proof: 26 + 52 = 78